# Towards Backdoor Stealthiness in Model Parameter Space

### v1

@@ 4.5. Backdoor Detection
Following previous works (Xu et al., 2024a, b), we choose five representative backdoor model detections for evaluation. The model detection refers to determining whether a given model is backdoored. We use 20 models for each poisoning rate with different random seeds. Then, we report the number of models detected as backdoor models out of the 20. Table 5 shows that all detections fall short when detecting Grond. In particular, NC (Wang et al., 2019), Tabor (Wang et al., 2019), and BTI-DBF (Xu et al., 2024a) can detect a small part of backdoored models, while FeatureRE (Wang et al., 2022a) and Unicorn cannot detect any of them. For featureRE (Wang et al., 2022a), we conjecture it is over-dependent on the separability in the feature space, but Grond does not rely on prominent backdoor features according to Figure 3. For Unicorn (Wang et al., 2023), the false positive rate is high, and it tends to report every class as the backdoor target, even on models trained with benign data only. Except for model detection, Grond can also bypass input-space detections as demonstrated in Appendix B.1.
4.6. Comparison with Supply-Chain Attacks
Supply-chain backdoor attacks assume the adversary could directly manipulate models’ parameters or control the backdoor training process for more powerful and stealthy backdoors. The backdoored model is provided as a service or final product to end users. Supply-chain backdoor attacks are attracting increasing industry and research community attention because of their stealthiness and significant real-world impact (Hong et al., 2022).
Sharing a similar threat model to supply-chain attacks, we compare Grond and three state-of-the-art supply-chain attacks, where these attacks are also designed to be robust against backdoor defenses. In particular, DFST (Cheng et al., 2021) proposes to include a controlled detoxification technique in the training process, which restrains the model from picking up simple features. DFBA (Cao et al., 2024) directly modifies a few parameters of a classifier to inject a backdoor. SSDT (Mo et al., 2024) introduces additional terms in the loss for the Source-Specific and Dynamic-Triggers (i.e., SSDT) attack, which obscures the difference between normal samples and malicious samples. Table 6 shows that supply-chain attacks can also be defeated by existing backdoor defenses. The ASR of DFST (Cheng et al., 2021), DFBA (Cao et al., 2024), and SSDT (Mo et al., 2024) are decreased to less than 10% while the BA drop is less than 3%.
Table 7. Comparsion with different strategies for the generation of backdoor triggers.
@@ 4.7. Ablation Study
Trigger generation. To explore the influence of trigger patterns, we employ and evaluate three types of triggers: random noise, PGD perturbation, and UPGD perturbation, using ResNet18 on CIFAR-10. The random noise is sampled from a uniform distribution, and the PGD employs a projected gradient descent to generate sample-wise perturbations (Madry et al., 2018). The generation of UPGD is described in Algorithm 1. All three triggers are limited to $8/255$ ($l_{\infty}$ norm) for imperceptibility and use the same training settings described in Table 11 in Appendix A.2.
@@
We show in Table 7 that random noise is ineffective as a backdoor trigger, with an ASR around 1%, even if no defense is applied. The sample-wise PGD perturbation is more effective than random noise and shows (limited) robustness against CLP and FT-SAM. UPGD generates the most effective backdoor trigger with an ASR higher than 80% after CLP and FT-SAM, and we speculate that the reason is that UPGD exploits features from the target class, similar to Narcissus (Zeng et al., 2023).
@@
Adversarial backdoor injection is critical. There are two components in Grond: the UPGD trigger generation and Adversarial Backdoor Injection. We conduct an ablation study with two architectures on CIFAR-10 to analyze the impact of the ABI component. As shown in Table 8, after removing the ABI component, CLP or FT-SAM can defend against the clean-label attack with the UPGD trigger. Thus, the Adversarial Backdoor Injection is the key component in maintaining the effectiveness of backdoor attacks against parameter-space defenses.
Dirty-label Grond. Grond works well in a clean-label setting that uses an invisible trigger and does not change the original labels of the poisoned samples. However, as we use the supply-chain threat model (the attacker has access to the training process), we could also explore the effect of a dirty-label backdoor attack. A dirty-label threat model could simplify the backdoor by poisoning samples from any class while the clean-label is limited to a single class.
As shown in Table 9, Grond with a dirty-label threat model can still perform well and even with a higher ASR than the clean-label setting. However, dirty-label Grond is less robust than clean-label Grond against backdoor defenses because it is less stealthy. For example, FT-SAM (Zhu et al., 2023) can decrease the ASR of dirty-label Grond (PR=5%) from 100% to 0%. When PR=1%, both CLP (Zheng et al., 2022) and FT-SAM (Zhu et al., 2023) can decrease the ASR from 100% to below 50%. We conjecture this is because the dirty-label setting obfuscates the benign semantics of images to their true labels, as the poisoned samples in each class have been assigned the backdoor target label. In contrast, in a clean-label setting, only the target class is poisoned.
Table 9. Evaluation of dirty-label Grond using ResNet18 on CIFAR-10.
Dirty-Label Grond No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
BA ASR BA ASR BA ASR
PR=5% 91.60 100 91.41 100 90.24 0.00
PR=1% 93.64 100 91.13 40.83 91.20 46.97
PR=0.5% 94.35 100 91.84 97.03 91.77 99.04
@@ 5. Stronger Defenders and Additional Analysis
This section dives deeper into stronger defenses and analysis of Grond. We evaluate Grond against a proactive defense in Section 5.1. Then, we provide analyses of Grond by Grad-CAM (defense) and feature spaces of different backdoor attacks in Section 5.2.
5.1. Proactivate Defense
Real-world powerful defenders could take more initiative by intervening proactively in the attack process and exploiting poisoned data. We evaluate Grond against the state-of-the-art proactive defense, CT (Qi et al., 2023b), that detects poisoned samples in the training data. Specifically, CT considers data from the original poisoned training data as regular batches and introduces randomly labeled benign data as confusing batches. Then, CT performs normal supervised training on both regular and confusing batches to produce an inference model, aiming to corrupt benign semantic features and correlations with correct labels in the inference model by confusing batches. The backdoor effect remains in the inference model because there is no trigger information in the confusing batches, and correctly predicted samples by the inference model are recorded as poisoned.
We apply CT to our adversarial backdoor injection process. Table 10 presents the detection results on two baseline attacks (BadNets (Gu et al., 2019) and Adap-patch (Qi et al., 2023a)) and Grond. CT is effective against the two baseline attacks, where most poisoned samples in the training set are detected with a relatively low false positive rate. However, CT is not capable of detecting Grond when the poisoning rate is lower than 5% due to a high false positive rate and low recall. To understand why CT is not effective against Grond, we recall the main idea of CT is to corrupt benign semantic features and their correct label but not corrupt backdoor semantic features. However, Grond utilizes the benign semantic features of the target class to generate UPGD perturbation as the trigger. CT’s mechanism also corrupts the backdoor features of Grond. Therefore, CT cannot effectively detects Grond poisoned samples.
@@ 5.2. Visualization
Grad-CAM cannot spot the trigger area of Grond. Grad-CAM (Selvaraju et al., 2017) was originally designed to visualize the network’s preference when taking an input image. In backdoor defense research, Grad-CAM is leveraged to highlight the important areas in order to detect the potential backdoor trigger area (Xu et al., 2020). Figure 7 shows the activated area of a clean model and Grond backdoored model using Grad-CAM. The activated area of Grond backdoored model is indistinguishable from the clean model, so the Grad-CAM-based defense (Xu et al., 2020) is also ineffective against Grond.
t-SNE visualization of feature space. Figure 6 shows the latent feature (feature space of the last convolutional layers) from Grond backdoor models with and without adversarial backdoor injection in 2-D space and other baselines attacks by t-SNE (Van der Maaten and Hinton, 2008). The poisoning rate for all is 0.5%. WaNet cannot achieve satisfactory ASR at this very low poisoning rate, so we use the default setting according to their open-source implementation. Specifically, we perform dimensionality reduction for the activation of the last convolutional layers by t-SNE. The model architecture is ResNet18 and trained on CIFAR-10. Each class of samples forms a tight cluster, and Grond poisoned samples are better mixed with the target class samples when the model is trained with adversarial backdoor injection.
@@ A.1. Datasets
CIFAR-10. The CIFAR-10 (Krizhevsky et al., 2009) contains 50,000 training images and 10,000 testing images with the size of $3\times 32\times 32$ in 10 classes.
@@ A.2. Backdoor Attacks
Like Narcissus, our attack uses the class bird (CIFAR10) as the target class. For ImageNet200, we use the stingray as the target class. The Grond poisoning rate (ImageNet200) used for results in Table 4 is 0.5%. For GTSRB, we use the speed limit (50) as the target class. The Grond poisoning rate (GTSRB) used for results in Table 4 is 1.74%. AdvDoor uses the same trigger and target class as ours. More details are provided in Table 11. For other attacks and hyperparameters not mentioned, we use the default setting from the original papers or open-source implementations.
A.3. Attack Summary
@@
A.4. Backdoor Defenses
NC [^14] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/neural_cleanse.py and Tabor [^15] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/tabor.py. We use the implementation from TrojanZoo (Pang et al., 2022). 1% training set and 100 epoch are used for trigger inversion.
A.5. Hyperparameters for Training Surrogate Models
Table 13. The settings for training surrogate models.
Config Value
Optimizer SGD, AdamW (InceptionNeXt)
Weight decay $5\times 10^{-4}$
learning rate 0.01 (CIFAR10, GTSRB), 0.001 (ImageNet200)
epoch 200 (GTSRB, CIFAR10), 100 (ImageNet200)
learning rate schedule MultiStepLR (100, 150) for CIFAR10 and GTSRB,
CosineAnnealingLR for ImageNet200
A.6. Hyperparameters for the Inversed Backdoor Feature Loss
Following the settings in BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask in Equation 4. The optimization of the mask uses 1% of training data. The $\lambda$ is 0.72. The elements in the mask are limited to continuous values between 0 and 1.
@@ A.6. Hyperparameters for the Inversed Backdoor Feature Loss
Table 15. Grond against defenses using different architectures on CIFAR-10 with a poisoning rate of 5%. The surrogate indicates the architecture used to generate UPGD as the trigger.
@@
Appendix B Additional Experiments
B.1. Detection of backdoor input.
@@ Appendix B Additional Experiments
B.2. Different Architectures with Different Surrogate Models
@@
We evaluate Grond with four additional victim architectures in Table 15: VGG16, DenseNet121, EfficienNet-B0, and InceptionNeXt-Tiny. In addition, as Grond requires a surrogate model to generate UPGD as the backdoor trigger, we provide the results when UPGD is generated using different architectures for the surrogate model. For each architecture, UPGD is generated by either the victim architecture or ResNet18 to perform our attack. In Table 15, we use the three most powerful defenses according to Tables 2 and 3. Regardless of the model’s architectures or the architectures for UPGD, Grond bypasses most defenses. This is because the UPGD contains semantic information of the target class and can be transferred among different architectures (Moosavi-Dezfooli et al., 2017). In a few cases, using UPGD generated by the same architecture shows better attack performance. For example, conducting Grond on InceptionNeXt-Tiny with UPGD generated by InceptionNeXt-Tiny shows ASRs above 90%, but also a much lower ASR when using UPGD generated by ResNet18. We conjecture that transferring UPGD from ResNet18 to InceptionNeXt-Tiny is more difficult than transferring it to other architectures due to the large convolution kernel design of InceptionNeXt.

### v2

@@ 4.6. Backdoor Detection
Following previous works (Xu et al., 2024a, b), we choose five representative backdoor model detections for evaluation. We use 20 models for each poisoning rate with different random seeds. Then, we report the number of models detected as backdoor models out of the 20. Table 6 shows that all detections fall short when detecting Grond. In particular, NC (Wang et al., 2019), Tabor (Wang et al., 2019), and BTI-DBF (Xu et al., 2024a) can detect a small part of backdoored models, while FeatureRE (Wang et al., 2022a) and Unicorn cannot detect any of them. For featureRE (Wang et al., 2022a), we conjecture it is over-dependent on the separability in the feature space, but Grond does not rely on prominent backdoor features according to Figure 9 in the Appendix App. C.4. For Unicorn (Wang et al., 2023), the false positive rate is high, and it tends to report every class as the backdoor target, even on models trained with benign data only. Except for model detection, Grond can also bypass input-space detections as demonstrated in App. C.1.
Table 7. Comparison with different strategies for the generation of backdoor triggers.
@@
Table 8. The semantic trigger (an automobile image) with different strategies for Grond. The “Target Same” refers to the target class being the same as the semantic trigger, i.e., automobile. The “Target Diff.” refers to the target class being different (i.e., airplane) from the semantic trigger.
Strategies for [Uncaptioned image] No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
BA ASR BA ASR BA ASR
Clean-Label Target Same 94.20 74.64 93.97 70.66 91.68 9.38
Target Diff. 93.82 94.42 93.75 93.50 90.92 31.37
Dirty-Label Target Same 93.77 100 93.88 100 91.40 12.17
Target Diff. 94.25 100 94.14 100 91.34 7.39
@@ 4.7. Ablation Study
The ablation study is designed for the two components of Grond, the UPGD trigger and Adversarial Backdoor Injection. In addition, we also evaluate the dirty-label setting of Grond and show the difference compared to using clean-label in App. C.5.
@@
Trigger generation. To explore the influence of trigger patterns, we employ and evaluate three types of triggers: random noise, PGD perturbation, and UPGD perturbation, using ResNet18 on CIFAR10. The random noise is sampled from a uniform distribution, and the PGD employs a projected gradient descent to generate sample-wise perturbations (Madry et al., 2018). The generation of UPGD is described in Algorithm 1. All three triggers are limited to $8/255$ ($l_{\infty}$ norm) for imperceptibility and use the same training settings described in Table 11 in App. A.2.
@@
Table 7 shows that random noise is ineffective as a backdoor trigger due to low ASR, even if no defense is applied. The sample-wise PGD perturbation is more effective than random noise and shows (limited) robustness against CLP and FT-SAM. UPGD generates the most effective backdoor trigger with an ASR higher than 80% after CLP and FT-SAM, and we speculate that the reason is that UPGD exploits features from the target class, similar to Narcissus (Zeng et al., 2023).
Concerning exploiting features from the target class, we also explore using the natural image as the trigger, which directly contains the semantic information. Table 8 shows the performance when using an automobile image from the CIFAR10 dataset as a semantic trigger. Specifically, inspired by naturally occurring backdoor (Khaddaj et al., 2023), we design the semantic trigger by resizing the automobile image to $8\times 8$ and sticking it on a part of (PR=5%) the training images. We consider four types of strategies, including only poisoning the same class as the trigger in clean-label, poisoning a different class (class airplane) in clean-label, only poisoning the same class as the trigger in dirty-label, and poisoning a different class (class airplane) in dirty-label. In Table 8, the semantic trigger can be effective as a backdoor trigger and resist against the CLP pruning. However, the semantic trigger is not robust against the FT-SAM fine-tuning. The reason is that the semantic trigger cannot effectively represent the feature of a class. Conversely, adversarial perturbation (PGD, UPGD) acquires more representative and discriminative class features, since they can capture the correlations among different regions of the decision boundary and easily fool most of the inputs (Moosavi-Dezfooli et al., 2017).
@@
Adversarial backdoor injection is critical. There are two components in Grond: the UPGD trigger generation and Adversarial Backdoor Injection. We conduct an ablation study with two architectures on CIFAR10 to analyze the impact of the ABI component. As shown in Table 9, after removing the ABI component, CLP or FT-SAM can defend against the clean-label attack with the UPGD trigger. Thus, the Adversarial Backdoor Injection is the key component in maintaining the effectiveness of backdoor attacks against parameter-space defenses.
@@ 5. Stronger Defenders and Additional Analysis
5.1. Proactive Defense
Real-world powerful defenders could take more initiative by intervening proactively in the attack process and exploiting poisoned data. We evaluate Grond against the SOTA proactive defense, CT (Qi et al., 2023b), that detects poisoned samples in the training data. Specifically, CT considers data from the original poisoned training data as regular batches and introduces randomly labeled benign data as confusing batches. Then, CT performs normal supervised training on both regular and confusing batches to produce an inference model, aiming to corrupt benign semantic features and correlations with correct labels in the inference model by confusing batches. The backdoor effect remains in the inference model because there is no trigger information in the confusing batches, and correctly predicted samples by the inference model are recorded as poisoned.
Table 10 presents the detection results on two baseline attacks (BadNets (Gu et al., 2019) and Adap-patch (Qi et al., 2023a)) and Grond. CT is effective against the two baseline attacks, where most poisoned samples in the training set are detected with a relatively low false positive rate. However, CT is not capable of detecting Grond when the poisoning rate is lower than 5% due to a high false positive rate and low recall. To understand why CT is not effective against Grond, we recall that the main idea of CT is to corrupt benign semantic features and their correct label, but not corrupt backdoor semantic features. However, Grond utilizes the benign semantic features of the target class to generate UPGD perturbation as the trigger. CT’s mechanism also corrupts the backdoor features of Grond. Therefore, CT cannot effectively detect Grond poisoned samples.
@@ 5.2. Visualization
Grad-CAM cannot spot the trigger area of Grond. Grad-CAM (Selvaraju et al., 2017) was originally designed to visualize the network’s preference when taking an input image. In backdoor defense research, Grad-CAM is leveraged to highlight the important areas in order to detect the potential backdoor trigger area (Cheng et al., 2020). Figure 8 shows the activated area of a clean model and Grond backdoored model using Grad-CAM. The activated area of Grond backdoored model is indistinguishable from the clean model, so the Grad-CAM-based defense (Cheng et al., 2020) is also ineffective against Grond.
t-SNE visualization of feature space. Figure 7 shows the latent feature (feature space of the last convolutional layers) from Grond backdoor models with and without adversarial backdoor injection in 2-D space and other baseline attacks by t-SNE (Van der Maaten and Hinton, 2008). The poisoning rate for all is 0.5%. WaNet cannot achieve satisfactory ASR at this very low poisoning rate, so we use the default setting according to their open-source implementation. Specifically, we perform dimensionality reduction for the latent features by t-SNE. The model architecture is ResNet18 and trained on CIFAR10. Each class of samples forms a tight cluster, and Grond poisoned samples are better mixed with the target class samples when the model is trained with adversarial backdoor injection.
@@ Appendix A Additional Details about Experimental Settings
Complete appendix and more experimental results can be found in the arXiv version of the paper.
@@ A.1. Datasets
CIFAR10. The CIFAR10 (Krizhevsky et al., 2009) contains 50,000 training images and 10,000 testing images with the size of $3\times 32\times 32$ in 10 classes.
@@ A.2. Backdoor Attacks
Like Narcissus, our attack uses the class bird (CIFAR10) as the target class. For ImageNet200, we use the “stingray” as the target class. The Grond poisoning rate (ImageNet200) used for results in Table 4 is 0.25%. For GTSRB, we use the speed limit (50) as the target class. The Grond poisoning rate (GTSRB) used for results in Table 4 is 1.74%. AdvDoor uses the same trigger and target class as ours. More details are provided in Table 11. For other attacks and hyperparameters not mentioned, we use the default setting from the original papers or open-source implementations.
A.3. Backdoor Defenses
NC [^13] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/neural_cleanse.py and Tabor [^14] https://github.com/ain-soph/trojanzoo/blob/main/trojanvision/defenses/backdoor/model_inspection/tabor.py. We use the implementation from TrojanZoo (Pang et al., 2022). 1% training set and 100 epochs are used for trigger inversion.
A.4. Hyperparameters for Training Surrogate Models
Table 12. The settings for training surrogate models.
Config Value
Optimizer SGD, AdamW (InceptionNeXt)
Weight decay $5\times 10^{-4}$
learning rate 0.01 (CIFAR10, GTSRB), 0.001 (ImageNet200)
epoch 200 (GTSRB, CIFAR10), 100 (ImageNet200)
learning rate schedule MultiStepLR (100, 150) for CIFAR10 and GTSRB,
CosineAnnealingLR for ImageNet200
Appendix B Attack Summary
@@
Table 15. Grond against defenses using different architectures on CIFAR10 with a poisoning rate of 5%. The surrogate indicates the architecture used to generate UPGD as the trigger.
@@
Appendix C Additional Experiments
C.1. Detection of backdoor input.
@@ Appendix C Additional Experiments
C.2. Different Architectures with Different Surrogate Models
@@
We evaluate Grond with four additional victim architectures in Table 15: VGG16, DenseNet121, EfficientNet-B0, and InceptionNeXt-Tiny. In addition, as Grond requires a surrogate model to generate UPGD as the backdoor trigger, we provide the results when UPGD is generated using different architectures for the surrogate model. For each architecture, UPGD is generated by either the victim architecture or ResNet18 to perform our attack. In Table 15, we use the three most powerful defenses according to Tables 2 and 3. Regardless of the model’s architectures or the architectures for UPGD, Grond bypasses most defenses. This is because the UPGD contains semantic information of the target class and can be transferred among different architectures (Moosavi-Dezfooli et al., 2017). In a few cases, using UPGD generated by the same architecture shows better attack performance. For example, conducting Grond on InceptionNeXt-Tiny with UPGD generated by InceptionNeXt-Tiny shows ASRs above 90%, but also a much lower ASR when using UPGD generated by ResNet18. We conjecture that transferring UPGD from ResNet18 to InceptionNeXt-Tiny is more difficult than transferring it to other architectures due to the large convolution kernel design of InceptionNeXt.
C.3. Large and Transformer-Based Architectures
To further evaluate the effectiveness of Grond on large-sized and especially transformer-based models, we conduct experiments on various architectures, including ResNet101 (He et al., 2016), Swin-Base (Liu et al., 2021), Swin-Large (Liu et al., 2021), ViT-Base (Dosovitskiy et al., 2021), and ViT-Large (Dosovitskiy et al., 2021).
There are two components (UPGD trigger and ABI) of Grond that are necessary to be applied to transformer-based architectures. The UPGD trigger generation algorithm is compatible with transformer-based architectures, as transformers’ gradient information can be used to generate UPGD perturbations. However, UCLC (Zheng et al., 2022) for ABI is incompatible with the attention mechanism. The UCLC is designed for feature channels of convolutional layers, which do not exist in the attention mechanism. Therefore, we design a tailored novel pruning method against backdoor neurons in the attention-based models. Specifically, we consider an attention head as a “neuron”, the new pruning method then calculates the singular value decomposition of each attention head as an index of relevance to the backdoor effect. The process can also be considered as solving for the approximation of the Lipschitz Constant for each attention head according to (Zheng et al., 2022). Then, the sensitive heads are updated using the following equation:
(4) ${\bm{\theta}}_{l}^{(k)}:=\left\{\begin{aligned} &{\bm{\theta}}_{l}^{(k)}\times 0.5,&\sigma({\bm{\theta}}_{l}^{(k)})>\text{mean}(\sigma({\bm{\theta}}_{l}))+u\times\text{std}(\sigma({\bm{\theta}}_{l}))\\ &{\bm{\theta}}_{l}^{(k)},&\text{otherwise,}\end{aligned}\right.$
where ${\bm{\theta}}_{l}^{(k)}$ is the weight of $k_{th}$ head at $l_{th}$ layer, $\sigma$ is the maximum of values in the singular value decomposition matrix of ${\bm{\theta}}_{l}^{(k)}$. We evaluate the performance of Grond on transformer architectures with three fine-tuning-based defenses. Table 16 provides the experimental results and demonstrates the effectiveness of Grond on large and transformer architectures. Notably, the three fine-tuning defenses cannot effectively decrease the ASR of Grond models. We conjecture this is because injecting a stealthy backdoor into larger models is easier.
The hyperparameters for training ResNet101 are the same as those in Table 11. For transformer models, the number of training epochs is 100 for both CIFAR10 and ImageNet200. The perturbation budget of UPGD for transformer models is $\epsilon=16$. The hyperparameters not mentioned are given in Table 11.
Table 16. The performance of Grond on large and transformer-based architectures.
Dataset Victim Surrogate Parameters (M) No Defense vanilla FT FT-SAM (Zhu et al., 2023) FST (Min et al., 2023)
BA ASR BA ASR BA ASR BA ASR
CIFAR10 ResNet101 ResNet18 42.51 93.75 99.22 87.37 99.92 92.78 99.04 93.65 99.77
ViT-Base ViT-Base 85.27 80.24 98.80 76.43 98.72 79.50 98.10 79.77 98.73
ViT-Large 302.60 78.92 98.89 74.47 97.53 77.98 98.56 78.53 98.53
Swin-Base Swin-Base 86.70 83.08 99.65 81.61 99.80 80.21 98.88 82.44 99.86
Swin-Large 194.93 83.04 99.69 80.79 99.35 80.99 99.47 79.31 99.53
ImageNet200 ViT-Base ViT-Base 85.95 83.41 98.36 82.08 98.86 83.48 98.24 80.99 97.84
ViT-Large 303.51 86.08 89.46 83.19 89.07 86.06 90.52 82.36 85.35
Swin-Base Swin-Base 86.95 74.68 97.82 72.56 98.29 75.75 96.11 72.29 94.77
Swin-Large 195.30 76.99 95.36 74.29 97.91 75.35 90.77 73.07 90.72
C.4. Backdoor analysis by feature space inversion
We also provide a feature space analysis for different attacks by using a feature mask to decouple the benign and backdoor features following BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b). The decoupling assumes that benign features related to the correct prediction introduce a lower loss, and the backdoor features related to backdoor prediction introduce a higher loss. In particular, the benign and inversed backdoor features are decoupled as follows:
where $g=f_{L-1}\circ\cdots f_{1}$ is the model without the classification head. ${\bm{m}}$ is the learned feature mask, and $\mathcal{D}_{l}$ is a small set of benign samples with correct labels. As validated with BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), benign and backdoor features can be decoupled by the mask ${\bm{m}}$, after which backdoor features will introduce a substantially higher loss with respect to the ground truth label.
Hyperparameters for the Inversed Backdoor Feature Loss Following the settings in BTI-DBF (Xu et al., 2024a) and BAN (Xu et al., 2024b), we use Adam and the learning rate of 0.01 to search for 20 epochs for the feature mask in Equation 5. The optimization of the mask uses 1% of the training data. $\lambda$ is 0.72. The elements in the mask are limited to continuous values between 0 and 1.
Table 17. Evaluation of dirty-label Grond using ResNet18 on CIFAR10. The dynamic target refers to samples from different classes having different backdoor targets.
Dirty-Label Grond PR No Defense CLP (Zheng et al., 2022) FT-SAM (Zhu et al., 2023)
BA ASR BA ASR BA ASR
One Target 5% 91.60 100 91.41 100 90.24 0.00
1% 93.64 100 91.13 40.83 91.20 46.97
0.5% 94.35 100 91.84 97.03 91.77 99.04
Dynamic Target 10% 94.26 91.95 94.13 77.29 91.67 1.46
5% 94.22 90.43 94.04 90.34 91.45 5.48
1% 94.33 81.82 94.26 81.80 92.36 15.77
C.5. Dirty-label Grond.
Grond works well in a clean-label setting that uses an invisible trigger and does not change the original labels of the poisoned samples. However, as we use the supply-chain threat model (the attacker has access to the training process), we could also explore the effect of a dirty-label backdoor attack. A dirty-label threat model could simplify the backdoor by poisoning samples from any class, while the clean-label is limited to a single class.
Table 17 considers two types of dirty-label Grond. The “One Target” uses the same setting and trigger as clean-label Grond, i.e., misclassifying any inputs with the trigger to the one target class. The “Dynamic Target” refers to misclassifying inputs from different classes to different targets. In addition, “Dynamic Target” Grond uses non-target UPGD as the trigger, which is generated by maximizing the loss between ${\bm{x}}+\bm{\delta}$ and the true label.
According to the results, One Target Grond achieves very high ASR. Dynamic Target Grond performs worse than One Target since Dynamic Target is a more complex task. Overall, dirty-label Grond can still perform well, and even has a higher ASR than the clean-label setting. However, dirty-label Grond is less robust than clean-label Grond against backdoor defenses because it is less stealthy. For example, FT-SAM can decrease the ASR of Dynamic Target Grond (PR=5%) from 94.22% to below 10%, and decrease One Target ASR from 100% to 0.0%. We conjecture this is because the dirty-label setting obfuscates the benign semantics of images to their true labels, as the poisoned samples in each class have been assigned the backdoor target label. In contrast, only the target class is poisoned in a clean-label setting.